for what value of x does
1 + 2 + 3+4+...+x=120?

Answer by rothauserc(4718)

(N(N+1)) / 2 = 120
multiply both sides of = by 2
N(N+1) = 240
N^2 +N -240 = 0
factor quadratic
(N+16)(N-15) = 0
N = -16 or N = 15
in this case, we want the positive value for N
N = x = 15
value of x is 15